OpenStudy (anonymous):

Please help me!!! :p Simplify 3^2 ⋅ 3^5

OpenStudy (kl0723):

they have the same base which is 3 so the base remains the same and then just add the exponents ;) so 3^2*3^5= same base which is 3^(2+5)=3^7

OpenStudy (anonymous):

awsome thanks!!!

OpenStudy (kl0723):

yw :)